. 4.    Channel Markups:  (always solved using formula  based on selling price – MU r )

.                 A manufacturer sells a $17.00 item to the wholesaler for $22.00.  The wholesaler in

.   turn sells it to the retailer for $28.00 and the retailer sells it to the public for $33.00.  

.               What is the markup at each level?  What does the consumer pay for the item?

Man. level = 23%            Wholesale level = 21%              Retail level = 15%

Consumer pays $33.00

So I know these answers are correct but I have no clue how to get the answers, please show step by step! thank you:)
Answer #1

Markup % = (Selling price - Cost price) / Cost price * 100

Manufacturer level

= (22 - 17) / 22*100

=22.7%

=23% (approximately)

Wholesaler level -

Markup % = (28 - 22) /28*100

=21.4

=21% ( approximately)

Retailer level -

Markup % = ( 33 - 28) / 33*100

=15.1%

=15% (approximately)

Retailers sell it to customer at $33. So, they will pay $33 for the product.

A Thumbs Up! Would be really helpful for me. If you have any questions, please leave a comment and I will get back to you as soon as possible.